The difference between two positive integers is 4. If the smaller is added to the square root of the larger, the sum is 86. find the integers.
Answer by mananth(16949)

let the integers be x x+4
..
x+sqrt(x+4)=86
sqrt(x+4)=86-x
square both sides
x+4=(86-x)^2
x+4 = 7396-172x+x^2
x^2-172x-x+7396-4=0
x^2-173x+7392=0
x^2-96x-77x+7392=0
x(x-96)-77(x-96)=0
(x-96)(x-77)=0
x= 96 OR 77
numbers are 77 & 81
..
m.ananth@hotmail.ca